# Dexterous Manipulation through Imitation Learning: A Survey

### v4

@@ III-E3 Discussion
BC Low Very Low (<1<1 ms) Low–Moderate (requires many demos; improves with online correction) Low Highly suitable for high-DOF real-time control; main bottleneck is data quality and covariate shift.
IRL High Low–Moderate (1–5 ms) Moderate–Low High Training is computationally expensive; inference feasible but may require reward computation if deployed online.
GAIL High Low–Moderate (2–5 ms) Moderate High Inference lightweight once trained; training unsuitable for strict real-time constraints.
HIL Moderate–High Moderate (2–10 ms) High (can reuse sub-policies) Moderate–High Two-tier policy execution increases latency; optimization needed for tight loops.
@@
III-F Comparative Analysis of Imitation Learning Approaches
@@ IV-A Two-fingered Traditional Gripper
Two-fingered grippers are widely used for their reliability, simplicity, and ease of control. Typically driven by a single actuator with one DoF, they are cost-effective and suitable for repetitive tasks requiring consistency [134]. For example, [135] demonstrated a Franka robot with such a gripper performing tasks like setting a breakfast table. Similarly, Kim et al. [136] used a two-fingered gripper for behavior cloning with gaze prediction, and a tendon-driven variant in [137] showed the ability to grasp diverse household objects.
Recent works have further extended gripper capabilities through large-scale imitation datasets such as MIME [138], RH20T [139], Bridge Data [140], and Droid [141]. Dual-arm systems also enhance manipulation by coordinating two grippers. For instance, [142] achieved banana peeling through dual-action IL. More complex tasks such as shrimp cooking, cloth folding, and dishwashing were demonstrated in Mobile ALOHA [143] and UMI [144].
Despite these advances, two-fingered grippers remain fundamentally limited in dexterous manipulation, which requires within-hand object reconfiguration [145]. Their simple structure and lack of internal DoFs restrict post-grasp adjustments [134]. Furthermore, morphological differences from the human hand hinder learning from demonstrations and prevent replication of human-like in-hand movements [146].
@@ IV-B Multi-fingered Anthropomorphic Hand
To overcome the dexterity limitations of two-fingered grippers, robotic hands with human-like morphology have been widely developed. These anthropomorphic hands are better suited for interacting with objects and environments designed for humans [147]. They can be typically classified by transmission mechanisms—tendon-driven, linkage-driven, direct-drive, and hybrid systems—which fundamentally affect their performance characteristics [148, 149].
Figure 3: Examples of multi-fingered anthropomorphic hands: (a) Shadow Dexterous Hand[150]; (b) Awiwi Hand [151]; (c) Biomimetic Hand [152]; (d) ILDA Hand [148]; (e) Hu et al.’s robotic hand [149]; (f) INSPIRE-ROBOTS RH56 Dexterous Hand[153]; (g) Linker Hand L20 [154]; (h) PUT-Hand [155]; (i) Allegro Hand[156]; (j) Faive Hand [157]; (k) Tesla Optimus Hand[158]; (l) Utah/MIT Dexterous Hand [159].
@@ IV-B1 Tendon-driven Approach
Representative examples include the Utah/MIT Hand [159] (see Fig. 3(l)), the Shadow Dexterous Hand [160] (see Fig. 3(a)), and the Awiwi Hand [161, 162, 163, 164] (see Fig. 3(b)), all of which adopt antagonistic tendon routing for biomimetic motion. The FLLEX Hand [165, 166] and Faive Hand [167] (see Fig. 3(j)) with rolling contact joints demonstrate robustness and ball-rolling manipulation, respectively. Other typical tendon-driven hands include the Robonaut R2 Hand [168], Valkyrie Hand [169, 170], UB Hand [171, 172, 173], DEXMART Hand [174, 175], iCub Hand [176], and Biomimetic Hand [177] (see Fig. 3(c)).
While remote actuation reduces hand weight, it introduces friction and tendon wear due to long transmission paths. To address this, some designs embed all actuators within the palm. Examples include the DEXHand [178], SpaceHand [179], CEA Hand [180], and OLYMPIC Hand [181], which prioritize modularity and compact integration. Commercial designs like DexHand 021 [182], Tesla Optimus Hand [183] (see Fig. 3(k)) and PUDU DH11 Hand [184] also follow this approach.
Despite their advantages in dexterity and anthropomorphism, tendon-driven hands face challenges such as friction loss [185, 186], end termination [163, 187, 188], tendon creep and wear [189, 190, 191], which impact durability and reliability. As a result, most remain within research settings, with limited deployment in real-world industrial applications.
@@ IV-B2 Linkage-driven Approach
Linkage-driven hands use rigid mechanical linkages to control joint motion, offering high precision, repeatability, and robustness. Compared to tendon-driven designs, they generally provide fewer DoFs but benefit from simpler, more reliable actuation. As a result, most commercial prosthetic and robotic hands adopt this mechanism. Due to space constraints and the demand for compactness, most linkage-driven fingers are actuated by a single motor and fall into two main categories: one-DoF coupled and multi-DoF underactuated types [192].
In the one-DoF type, joints are mechanically coupled, so preshaping remains fixed during flexion. Typical designs include the S-finger with inverse four-bar coupling [193] and the humanoid hand by Liu et al. using two four-bar linkages per finger [194]. Similar configurations appear in hands like the INSPIRE-ROBOTS RH56 [153] (see Fig. 3(f)), Bebionic Hand [195, 196], BrainRobotics Hand [197] and OYMotion OHand [198].
In contrast, underactuated fingers can adapt to contact forces, enhancing grasp adaptability. Examples include the Southampton Hand with a Whiffle tree mechanism [199], the LISA Hand using linkage-based self-adaptation [200]; MORA HAP-2 Hand [201], AR Hand III [202], and Cheng et al.’s prosthetic hand [203] with multi-bar or four-bar adaptive linkages.
While most designs use one motor per finger, a few incorporate multiple actuators for higher dexterity. The ILDA Hand [148] (see Fig. 3(d)) employs three motors per finger with combined PSS/PSU chains and four-bar linkages, achieving workspace and fingertip force comparable to human hands. Similar high-DoF linkage designs appear in the Linker Hand L20 [154](see Fig. 3(g)), AIDIN ROBOTICS Hand [204] and the RY-H1 Hand [205].
@@ IV-B3 Direct-driven Approach
Direct-drive hands eliminate intermediate transmission by connecting actuators directly to joints. This simplifies mechanical structure while still allowing for high actuatable DoFs, similar to tendon-driven designs.
Representative examples include the OCU-Hand [206] with 19 DoFs, where most joints are individually driven by embedded DC motors, and the TWENDY-ONE hand [207], which achieves 13 DoFs via joint-level motor placement. The KITECH-Hand [208], Allegro Hand [156] (see Fig. 3(i)) and LEAP Hand [209] adopt modular finger designs, integrating motors directly into the phalanges. The LEAP Hand also introduces a novel universal abduction-adduction motor configuration for enhanced MCP joint flexibility.
@@ IV-B4 Hybrid-transmission Approach
For example, the DLR/HIT Hand II [210] and NAIST Hand [211] use modular fingers with a combination of motors, belts, gears, tendon or linkage systems. The MCR-Hand series [212, 213] utilize a linkage-tendon mixed transmission system to achieve compactness and high functionality. Adab Mora Hand [214], LEAP Hand V2 (DLA Hand) [215] and Hu et al.’s hand [149] (see Fig. 3(e)) also integrate multiple transmission elements within fingers to enhance overall performance and adaptability.
Other hybrid designs explicitly differentiate mechanisms across fingers to match specific functional needs. The PUT-Hand [155] (see Fig. 3(h)), for instance, combines a direct-drive thumb, linkage-driven index/middle fingers, and tendon-driven ring/little fingers. Similarly, the MPL v2.0 Hand [216], Tact Hand [217], and Six-DoF Open Source Hand [218] adopt direct or geared actuation for the thumb while using tendons, linkages or timing belts for other fingers. Hands developed by Owen et al. [219], Ryu et al. [220], and Ke et al. [221] follow a similar approach, implementing thumb-specific hybrid strategies to enhance opposability and dexterity.
@@
Shadow Dexterous Hand[150] 20 / 24 Remote tendon-driven (20 Smart Motors in forearm) Fully biomimetic; 129 sensors; $\pm$1° joint precision High-dexterity research, tele-operation, AI/ML benchmarking
Allegro Hand[156] 16 / 16 Direct-driven (Maxon DC motors in each phalanx) Modular finger units; ROS-native; low backlash Academic RL/IL experiments, sim-to-real transfer
Linker Hand (L20)[154] 16 / 21 Linkage-driven (4 motors per finger) ROS-native; piezoresistive sensors; optional visual-tactile perception High-dexterity research, Industrial applications
BarrettHand (BH8-280)[222] 4 / 7 Linkage-driven (1 motor per finger + spread) Under-actuated adaptive fingers; rugged mechanics Industrial grasping, pick-and-place, educational labs
LEAP Hand[209] 16 / 16 Direct-driven (compact joint motors) Low-cost, open-source, anthropomorphic yet lightweight Learning-focused projects, mobile manipulation
i-HY Hand (3-finger)[223] 9 / 9 Tendon-driven (elastic tendons + capstan) Compliant, under-actuated; high-impact robustness Robust in-hand manipulation, field robotics
@@ IV-C Three-fingered Robotic Claw: A Trade-off Solution
The diversity of anthropomorphic hand designs largely stems from a fundamental trade-off between mechanical simplicity and dexterous capability [224]. While the human hand has over 20 DoFs [225, 226, 227, 228, 229, 230, 231], replicating this complexity mechanically remains impractical. Higher dexterity often increases structural and control complexity, cost, and susceptibility to failure [223, 232, 233], limiting the feasibility of high-DoF hands in real-world applications [234, 235].
To mitigate these issues, several simplification strategies are adopted: underactuation with elastic components [232, 233, 150, 158, 199, 200], reducing non-essential DoFs or phalanges [158, 153, 195, 198], or even omitting a finger entirely [159, 169, 178, 207, 209, 156]. These approaches highlight the challenge of maximizing functionality within practical constraints.
Figure 4: Examples of three-fingered robotic claws: (a) DEX-$E$ [236]; (b) BarrettHand [222]; (c) i-HY Hand [223]; (d) DoraHand [237].
As a compromise between the minimalist two-fingered gripper and complex multi-fingered anthropomorphic hands, the three-fingered robotic claw offers a functional middle ground. Though not anatomically human-like, three fingers are sufficient for executing common grasp types such as cylindrical and spherical power grasps [223], and can support a subset of in-hand manipulation tasks.
Numerous three-fingered claws have demonstrated impressive capabilities. For example, Shadow’s DEX-EE [236] (see Fig. 4(a)) and the TRX Hand [238] exhibit high robustness and dexterity. The BarrettHand [239] (see Fig. 4(b)) achieves adaptive grasping through underactuation. Tendon-driven designs like the i-HY Hand [223] (see Fig. 4(c)) and Model O [240] enable actions such as pivoting and precision transitions. Systems such as DClaw [241] and TriFinger [242] are capable of performing fine manipulation tasks via RL. Other novel architectures include linkage-based [243], motor-multiplexed [244], and link-belt-integrated claws [245], each offering different characteristics.
Three-fingered claws such as the DoraHand [235] (see Fig. 4(d)), SARAH [246], D’Manus [247], and Kinova Jaco’s claw [248] further demonstrate the practicality and versatility of this design choice in both research and assistive applications. Based on the foregoing analysis, the key characteristics of representative robotic hands across the three end-effector categories are summarized in Table V.

### v5

@@ III-E3 Discussion
BC Low Very Low (<1<1 ms) Low–Moderate (requires many demos; improves with online correction) Low Highly suitable for high-DOF real-time control; the main bottleneck is data quality and covariate shift.
IRL High Low–Moderate (1–5 ms) Moderate–Low High Training is computationally expensive; inference is feasible but may require reward computation if deployed online.
GAIL High Low–Moderate (2–5 ms) Moderate High Inference is lightweight once trained; training is unsuitable for strict real-time constraints.
HIL Moderate–High Moderate (2–10 ms) High (can reuse sub-policies) Moderate–High Two-tier policy execution increases latency; optimization is needed for tight loops.
@@ III-F Comparative Analysis of Imitation Learning Approaches
In addition to these strengths and weaknesses, ongoing debates highlight critical trade-offs that influence their application in complex dexterous manipulation tasks. (1) A central debate revolves around the balance between simplicity and scalability. BC is simple and data-efficient, making it effective for tasks with lower complexity or shorter horizons, but it struggles with generalization in high-DoF dexterous manipulation tasks. In contrast, IRL and GAIL offer better generalization, with IRL excelling at capturing expert behavior for unseen states but facing scalability issues due to the challenge of reward design. GAIL, on the other hand, provides greater scalability by eliminating the need for explicit reward functions, but it is prone to instability and mode collapse, particularly in high-DoF or multi-modal tasks. (2) Another ongoing discussion centers on the trade-off between end-to-end learning from raw observations and using intermediate representations. End-to-end learning simplifies the pipeline by directly mapping raw inputs to outputs, but it often requires large amounts of data and struggles with generalization, especially in complex tasks. In contrast, using intermediate representations, such as semantic keypoints, can improve efficiency and generalization by focusing on task-relevant abstractions. However, this approach introduces added complexity in terms of feature extraction and domain-specific design, increasing the computational burden. (3) A further debate concerns the trade-off between learning from expert demonstrations versus self-exploration. Learning from expert demonstrations (e.g., via BC or GAIL) can offer highly effective solutions by mimicking human-like actions, but it relies heavily on high-quality expert data, which may be scarce or difficult to obtain. Self-exploration, however, allows the model to learn through trial and error, leading to greater flexibility in task environments. Yet, this approach often suffers from inefficiencies and a longer learning time, as it requires the agent to explore numerous states to achieve a sufficient understanding of task-specific behaviors.
@@ IV-A Two-fingered Traditional Gripper
Two-fingered grippers are widely used for their reliability, simplicity, and ease of control. Typically driven by a single actuator with one DoF, they are cost-effective and suitable for repetitive tasks requiring consistency [yu_dexterous_2022]. For example, [nasiriany2022learningretrievalpriordata] demonstrated a Franka robot with such a gripper performing tasks like setting a breakfast table. Similarly, Kim et al. [kim2020using] used a two-fingered gripper for behavior cloning with gaze prediction, and a tendon-driven variant in [ciocarlie_kinetic_2013] showed the ability to grasp diverse household objects.
Recent works have further extended gripper capabilities through large-scale imitation datasets such as MIME [sharma_multiple_2018], RH20T [fang_rh20t_2024], Bridge Data [ebert_bridge_2022], and Droid [khazatsky_droid_2024]. Dual-arm systems also enhanced manipulation by coordinating two grippers. For instance, [kim_goal-conditioned_2024] achieved banana peeling through dual-action IL. More complex tasks such as shrimp cooking, cloth folding, and dishwashing were demonstrated in Mobile ALOHA [fu2024mobile] and UMI [chi2024universal].
Despite these advances, two-fingered grippers remain fundamentally limited in dexterous manipulation, which requires within-hand object reconfiguration [ma_dexterity_2011]. Their simple structure and lack of internal DoFs restrict post-grasp adjustments [yu_dexterous_2022]. Furthermore, morphological differences from the human hand hinder learning from demonstrations and prevent replication of human-like in-hand movements [kadalagere_sampath_review_2023].
@@ IV-B Multi-fingered Anthropomorphic Hand
To overcome the dexterity limitations of two-fingered grippers, robotic hands with human-like morphology have been widely developed. These anthropomorphic hands are better suited for interacting with objects and environments designed for humans [vianello_human-humanoid_2021]. They can be typically classified by transmission mechanisms—tendon-driven, linkage-driven, direct-drive, and hybrid systems—which fundamentally affect their performance characteristics [kim_integrated_2021, hu_design_2023].
Figure 3: Examples of multi-fingered anthropomorphic hands: (a) Shadow Dexterous Hand[shadow_hand_web2]; (b) Awiwi Hand [Awiwi_Hand_web]; (c) Biomimetic Hand [Biomimetic_Hand_web]; (d) ILDA Hand [kim_integrated_2021]; (e) Hu et al.’s robotic hand [hu_design_2023]; (f) INSPIRE-ROBOTS RH56 Dexterous Hand[inspire_RH56hand_web]; (g) Linker Hand L20 [linkerhand]; (h) PUT-Hand [mankowski_put-handhybrid_2020]; (i) Allegro Hand[Allegro__Hand_web]; (j) Faive Hand [Faive_hand_web]; (k) Tesla Optimus Hand[Tesla_Optimus_web]; (l) Utah/MIT Dexterous Hand [jacobsen_design_1986].
@@ IV-B1 Tendon-driven Approach
Representative examples include the Utah/MIT Hand [jacobsen_design_1986] (see Fig. 3(l)), the Shadow Dexterous Hand [shadow_hand_web] (see Fig. 3(a)), and the Awiwi Hand [grebenstein_dlr_2011, grebenstein_hand_2012, grebenstein_antagonistically_2010, grebenstein_antagonism_2008] (see Fig. 3(b)), all of which adopt antagonistic tendon routing for biomimetic motion. The FLLEX Hand [kim_fluid_2019, FLLEX_Hand_v2_web] and Faive Hand [toshimitsu_getting_2023] (see Fig. 3(j)) with rolling contact joints demonstrate robustness and ball-rolling manipulation, respectively. Other typical tendon-driven hands include the Robonaut R2 Hand [bridgwater_robonaut_2012], Valkyrie Hand [radford_valkyrie_2015, guo_design_2014], UB Hand [lotti_development_2005, palli_development_2012, melchiorri_development_2013], DEXMART Hand [palli_dexmart_2014, siciliano_advanced_2012], iCub Hand [sureshbabu_new_2015], and Biomimetic Hand [zhe_xu_design_2016] (see Fig. 3(c)).
While remote actuation reduces hand weight, it introduces friction and tendon wear due to long transmission paths. To address this, some designs embed all actuators within the palm. Examples include the DEXHand [chalon_dexhand_2011], SpaceHand [chalon_spacehand_2015], CEA Hand [martin_design_2014], and OLYMPIC Hand [liow_olympic_2020], which prioritize modularity and compact integration. Commercial designs like DexHand 021 [DexRobot_DexHand021_web], Tesla Optimus Hand [leddy_underactuated_2024] (see Fig. 3(k)) and PUDU DH11 Hand [PUDU_DH11_hand_web] also follow this approach.
Despite their advantages in dexterity and anthropomorphism, tendon-driven hands face challenges such as friction loss [reinecke_guiding_2014, uchiyama_method_1995], end termination [grebenstein_antagonistically_2010, grebenstein_approaching_2014, gerez_compact_2018], tendon creep and wear [palli_modeling_2012, friedl_frcef_2015, goswami_dlr_2017], which impact durability and reliability. As a result, most remain within research settings, with limited deployment in real-world industrial applications.
@@ IV-B2 Linkage-driven Approach
Linkage-driven hands use rigid mechanical linkages to control joint motion, offering high precision, repeatability, and robustness. Compared to tendon-driven designs, they generally provide fewer DoFs but benefit from simpler, more reliable actuation. As a result, most commercial prosthetic and robotic hands adopt this mechanism. Due to space constraints and the demand for compactness, most linkage-driven fingers are actuated by a single motor and fall into two main categories: one-DoF coupled and multi-DoF underactuated types [kashef_robotic_2020].
In the one-DoF type, joints are mechanically coupled, so preshaping remains fixed during flexion. Typical designs include the S-finger with inverse four-bar coupling [imbinto_s-finger_2018] and the humanoid hand by Liu et al. using two four-bar linkages per finger [liu_development_2016]. Similar configurations appear in hands like the INSPIRE-ROBOTS RH56 [inspire_RH56hand_web] (see Fig. 3(f)), Bebionic Hand [bebionic_Hand_web, medynski_bebionic_2011], BrainRobotics Hand [brainrobotics_Hand_web] and OYMotion OHand [OYMotion_OHand_web].
In contrast, underactuated fingers can adapt to contact forces, enhancing grasp adaptability. Examples include the Southampton Hand with a Whiffle tree mechanism [kyberd_design_2001], the LISA Hand using linkage-based self-adaptation [jin_lisa_2012]; MORA HAP-2 Hand [gopura_prosthetic_2017], AR Hand III [yang_anthropomorphic_2009], and Cheng et al.’s prosthetic hand [cheng_design_2017] with multi-bar or four-bar adaptive linkages.
While most designs use one motor per finger, a few incorporate multiple actuators for higher dexterity. The ILDA Hand [kim_integrated_2021] (see Fig. 3(d)) employs three motors per finger with combined PSS/PSU chains and four-bar linkages, achieving workspace and fingertip force comparable to human hands. Similar high-DoF linkage designs appear in the Linker Hand L20 [linkerhand](see Fig. 3(g)), AIDIN ROBOTICS Hand [aidinRobotics_Hand_web] and the RY-H1 Hand [Ruiyan_RYH1_Hand_web].
@@ IV-B3 Direct-driven Approach
Direct-drive hands eliminate intermediate transmission by connecting actuators directly to joints. This simplifies the mechanical structure while still allowing for high actuatable DoFs, similar to tendon-driven designs.
Representative examples include the OCU-Hand [mahmoud_dexterous_2010] with 19 DoFs, where most joints are individually driven by embedded DC motors, and the TWENDY-ONE hand [iwata_design_2009], which achieves 13 DoFs via joint-level motor placement. The KITECH-Hand [lee_kitech-hand_2017], Allegro Hand [Allegro__Hand_web] (see Fig. 3(i)) and LEAP Hand [shaw_leap_2023] adopt modular finger designs, integrating motors directly into the phalanges. The LEAP Hand also introduces a novel universal abduction-adduction motor configuration for enhanced MCP joint flexibility.
@@ IV-B4 Hybrid-transmission Approach
For example, the DLR/HIT Hand II [liu_multisensory_2008] and NAIST Hand [ueda_development_2005] use modular fingers with a combination of motors, belts, gears, tendons, or linkage systems. The MCR-Hand series [yang_affordable_2021, yang_low-cost_2021] utilizes a linkage-tendon mixed transmission system to achieve compactness and high functionality. Adab Mora Hand [abayasiri_under-actuated_2020], LEAP Hand V2 (DLA Hand) [shaw_leap_2024], and Hu et al.’s hand [hu_design_2023] (see Fig. 3(e)) also integrate multiple transmission elements within fingers to enhance overall performance and adaptability.
Other hybrid designs explicitly differentiate mechanisms across fingers to match specific functional needs. The PUT-Hand [mankowski_put-handhybrid_2020] (see Fig. 3(h)), for instance, combines a direct-drive thumb, linkage-driven index/middle fingers, and tendon-driven ring/little fingers. Similarly, the MPL v2.0 Hand [johannes_overview_2011], Tact Hand [slade_tact_2015], and Six-DoF Open Source Hand [krausz_design_2016] adopt direct or geared actuation for the thumb while using tendons, linkages or timing belts for other fingers. Hands developed by Owen et al. [owen_development_2018], Ryu et al. [ryu_development_2020], and Ke et al. [ke_new_2021] follow a similar approach, implementing thumb-specific hybrid strategies to enhance opposability and dexterity.
@@
Shadow Dexterous Hand[shadow_hand_web2] 20 / 24 Remote tendon-driven (20 Smart Motors in forearm) Fully biomimetic; 129 sensors; $\pm$1° joint precision High-dexterity research, tele-operation, AI/ML benchmarking
Allegro Hand[Allegro__Hand_web] 16 / 16 Direct-driven (Maxon DC motors in each phalanx) Modular finger units; ROS-native; low backlash Academic RL/IL experiments, sim-to-real transfer
Linker Hand (L20)[linkerhand] 16 / 21 Linkage-driven (4 motors per finger) ROS-native; piezoresistive sensors; optional visual-tactile perception High-dexterity research, Industrial applications
BarrettHand (BH8-280)[BarrettHand_web] 4 / 7 Linkage-driven (1 motor per finger + spread) Under-actuated adaptive fingers; rugged mechanics Industrial grasping, pick-and-place, educational labs
LEAP Hand[shaw_leap_2023] 16 / 16 Direct-driven (compact joint motors) Low-cost, open-source, anthropomorphic yet lightweight Learning-focused projects, mobile manipulation
i-HY Hand (3-finger)[odhner_compliant_2014] 9 / 9 Tendon-driven (elastic tendons + capstan) Compliant, under-actuated; high-impact robustness Robust in-hand manipulation, field robotics
@@ IV-C Three-fingered Robotic Claw: A Trade-off Solution
The diversity of anthropomorphic hand designs largely stems from a fundamental trade-off between mechanical simplicity and dexterous capability [bicchi_hands_2000]. While the human hand has over 20 DoFs [pena_pitarch_virtual_2008, yang_development_2008, samadani_multi-constrained_2012, lenarcic_robot_2013, prabaharan_nadine_2017, zhou_bcl-13_2018, zarzoura_investigation_2019], replicating this complexity mechanically remains impractical. Higher dexterity often increases structural and control complexity, cost, and susceptibility to failure [odhner_compliant_2014, demers_kinematic_2011, mnyusiwalla_new_2016], limiting the feasibility of high-DoF hands in real-world applications [feix_metric_2013, wang_dorahand_2022].
To mitigate these issues, several simplification strategies are adopted: underactuation with elastic components [demers_kinematic_2011, mnyusiwalla_new_2016, shadow_hand_web2, Tesla_Optimus_web, kyberd_design_2001, jin_lisa_2012], reducing non-essential DoFs or phalanges [Tesla_Optimus_web, inspire_RH56hand_web, bebionic_Hand_web, OYMotion_OHand_web], or even omitting a finger entirely [jacobsen_design_1986, radford_valkyrie_2015, chalon_dexhand_2011, iwata_design_2009, shaw_leap_2023, Allegro__Hand_web]. These approaches highlight the challenge of maximizing functionality within practical constraints.
Figure 4: Examples of three-fingered robotic claws: (a) DEX-$E$ [Shadow_DEXEE_web]; (b) BarrettHand [BarrettHand_web]; (c) i-HY Hand [odhner_compliant_2014]; (d) DoraHand [DoraHand_web].
As a compromise between the minimalist two-fingered gripper and complex multi-fingered anthropomorphic hands, the three-fingered robotic claw offers a functional middle ground. Though not anatomically human-like, three fingers are sufficient for executing common grasp types such as cylindrical and spherical power grasps [odhner_compliant_2014], and can support a subset of in-hand manipulation tasks.
Numerous three-fingered claws have demonstrated impressive capabilities. For example, Shadow’s DEX-EE [Shadow_DEXEE_web] (see Fig. 4(a)) and the TRX Hand [hu_dexterous_2023] exhibited high robustness and dexterity. The BarrettHand [townsend_barretthand_2000] (see Fig. 4(b)) achieved adaptive grasping through underactuation. Tendon-driven designs like the i-HY Hand [odhner_compliant_2014] (see Fig. 4(c)) and Model O [ma_yale_2017] enabled actions such as pivoting and precision transitions. Systems such as DClaw [zhu_dexterous_2019] and TriFinger [wuthrich_trifinger_2021] were capable of performing fine manipulation tasks via RL. Other novel architectures include linkage-based [li_linkage-driven_2024], motor-multiplexed [xu_muxhand_2024], and link-belt-integrated claws [kim_blt_2020], each offering different characteristics.
Three-fingered claws such as the DoraHand [wang_dorahand_2022] (see Fig. 4(d)), SARAH [laliberte_underactuation_2001], D’Manus [bhirangi_all_2024], and Kinova Jaco’s claw [Kinova_arm_web] further demonstrated the practicality and versatility of this design choice in both research and assistive applications. Based on the foregoing analysis, the key characteristics of representative robotic hands across the three end-effector categories are summarized in Table V.